these r my questions i need the answers 4 tomorrow so pls figure them out
u gotta figure out the value of "x"
7x=7=5x-11 3x=13=9-5x 4x-8=6x-14

Answer by checkley71(8403)

IN THE FIRST 2 EQUATIONS YOU HAVE ONE TOO MANY = SIGNS.
CORRECT & RESUBMIT THESE 2.
4X-8=6X-14
4X-6X=-14+8
-2X=-6
X=-6/-2
X=3 ANSWER.
PROOF:
4*3-8=6*3-14
12-8=18-14
4=4